# Learning Mamba as a Continual Learner

### v1

@@ Abstract
Continual learning (CL) aims to efficiently learn and accumulate knowledge from a data stream with different distributions. By formulating CL as a sequence prediction task, meta-continual learning (MCL) enables to meta-learn an efficient continual learner based on the recent advanced sequence models, e.g., Transformers. Although attention-free models (e.g., Linear Transformers) can ideally match CL’s essential objective and efficiency requirements, they usually perform not well in MCL. Considering that the attention-free Mamba achieves excellent performances matching Transformers’ on general sequence modeling tasks, in this paper, we aim to answer a question – Can attention-free Mamba perform well on MCL? By formulating Mamba with a selective state space model (SSM) for MCL tasks, we propose to meta-learn Mamba as a continual learner, referred to as MambaCL. By incorporating a selectivity regularization, we can effectively train MambaCL. Through comprehensive experiments across various CL tasks, we also explore how Mamba and other models perform in different MCL scenarios. Our experiments and analyses highlight the promising performance and generalization capabilities of Mamba in MCL.
@@ 1 Introduction
Continual learning (CL) aims to efficiently learn and accumulate knowledge in a non-stationary data stream (De Lange et al., 2021; Wang et al., 2024) containing different tasks. Given a sequence of data ${\mathcal{D}}_{T}=(({\mathbf{x}}_{1},y_{1}),...,({\mathbf{x}}_{t},y_{t}),...,({\mathbf{x}}_{T},y_{T}))$ with a series of paired observations ${\mathbf{x}}_{i}$ (e.g., images) and targets $y_{i}$ (e.g., class labels) from different tasks, CL is usually produced to learn one model $P_{\phi_{t}}(y|{\mathbf{x}})$ parameterized by $\phi_{t}$ that can perform prediction for any tasks corresponding to the seen data ${\mathcal{D}}_{t}$. For example, in class incremental learning (CIL) (Rebuffi et al., 2017; Zhou et al., 2023), a widely studied CL scenario, ${\mathcal{D}}_{T}$ consists of data with incrementally added classes, and $P_{\phi_{t}}(y|{\mathbf{x}})$ is trained to recognize all previously seen classes. To ensure computational and memory efficiency, CL methods are explored for learning from data streams while minimizing the storage of historical data or limiting running memory growth, such as restricting the increase rate to be constant or sub-linear (De Lange et al., 2021; Ostapenko et al., 2021). The main challenge in CL is to preserve performance on previously seen tasks while continually updating the model parameters $\phi_{t}$ (De Lange et al., 2021; Wang et al., 2024).
@@
CL methods continually train/update the model $P_{\phi_{t}}(y|{\mathbf{x}})$ from seen sequence ${\mathcal{D}}_{t}$ at arbitrary step $t$ and perform predictions on any observation ${\mathbf{x}}^{\text{test}}$ (following seen data distribution) for the corresponding $y^{\text{test}}$. From this perspective, the whole learning and inference process in CL can be seen as a sequence prediction (SP) problem, i.e., predicting $y_{\text{test}}$ of a query ${\mathbf{x}}_{\text{test}}$ conditioning on the seen data sequence and the testing input, i.e., $({\mathcal{D}}_{t}^{\text{train}},{\mathbf{x}}^{\text{test}})\equiv({\mathbf{x}}_{1}^{\text{train}},y_{1}^{\text{train}},...,{\mathbf{x}}_{t}^{\text{train}},y_{t}^{\text{train}},{\mathbf{x}}^{\text{test}})$ (Lee et al., 2023; Bornschein et al., 2024). In conventional CL, the model parameter $\phi_{t}$ is trained to maintain the states on the sequence, i.e., knowledge in the historical data, in a way of $\phi_{t+1}=\texttt{optim-step}(\phi_{t},{\mathbf{x}}_{t},y_{t})$. This connection between sequence prediction and CL training process motivates us to investigate meta-learning a continual learner as a sequence prediction model, for computation-and-data-efficient CL. Through meta-continual learning (MCL) framework (Lee et al., 2023; Son et al., 2023), a continual learner $f_{\theta}()$ parameterized by $\theta$ is trained via sequence prediction on multiple CL episodes. A meta-learned $f_{\theta}()$ can take a given sequence $({\mathcal{D}}_{t},{\mathbf{x}}_{t+1})$ as input and predict the label $y_{t+1}=f_{\theta}(({\mathcal{D}}_{t},{\mathbf{x}}_{t+1}))$, which is equivalent to a predictive model conditioning on the seen data stream $P_{\theta}(y|{\mathbf{x}},{\mathcal{D}}_{t})$. The data stream can also be seen as a context of the tasks for performing prediction for a new query.
Transformers (Vaswani et al., 2017; Touvron et al., 2023) have shown strong sequence modeling capabilities and next-token prediction performance in language modeling (LM), relying on self-attention across per-step tokens and emergent in-context learning (ICL) ability (Brown et al., 2020; Garg et al., 2022). It is thus straightforward to meta-learn a Transformer as the SP-based continual learner (Lee et al., 2023; Son et al., 2023). Given a data stream in CL, a meta-learned Transformer generates a new key-value pair at each step and makes the prediction for each query based on attention over the key-value pairs retained from all preceding training samples. Benefiting from the retrieval-based modeling, Transformers can perform effectively in continual learning (CL) (Lee et al., 2023; Bornschein et al., 2024). However, they require maintaining key-value pairs for all seen training samples in a key-value cache, allowing the model to access all seen samples during inference. It contradicts the principles and intended purpose of continual learning. Although the key-value cache can be viewed as the hidden state of a recurrent neural network (RNN) (Katharopoulos et al., 2020; Lee et al., 2023), analogous to the parameters of a learner, its size grows linearly with the number of all seen tokens and suffers from increasing memory and computational demands over time. Despite their advanced sequence modeling capabilities as in (Lee et al., 2023), Transformers may not be an ideal choice for continual learning due to misalignment with the objectives of CL and efficiency concerns. A series of attention-free models achieve efficiency by approximating the softmax attention with kernel methods and linear operations, leading to constant hidden state sizes and linear computation complexity, such as Linear Transformer (Katharopoulos et al., 2020) and Performer (Choromanski et al., 2020). Although these efficient Transformers align better with the purpose of CL, it is seen that they cannot perform well in MCL (Lee et al., 2023), due to limitations in approximation and insufficient expressive power (Katharopoulos et al., 2020; Choromanski et al., 2020; Tay et al., 2020).
@@
Recent advancements of the state space models (SSMs) on sequence modeling lead to a series of attention-free models that are efficient in processing long sequences with nearly linear computation (Gu et al., 2021a; b). By integrating time-varying modeling into the SSM as a selective SSM, Mamba (Gu & Dao, 2023; Dao & Gu, 2024) can achieve near state-of-the-art performances on sequence modeling tasks (e.g., LM tasks (Gao et al., 2020)). Given its exceptional performance as an attention-free model with a constant hidden state size, which ideally aligns with the requirements of MCL, rather than relying on Transformers (Lee et al., 2023), we pose a concrete question: Can the attention-free model Mamba perform well in MCL? In this paper, we investigate this question by formulating the selective SSM and Mamba to handle MCL, referred to as MambaCL. We identify that it is not trivial to train the sequence prediction models, including Mamba, for MCL, due to difficulty in convergence. To address the issue, we introduce a selectivity regularizer relying on the connection across SSM/Mamba and Linear Transformers and Transformers, which guides the behaviour of the generated time-variant parameters of the selective SSM during training. Relying on the specifically designed regularization and customized designs, we achieve an effective MambaCL model for MCL. Beyond the scope of the existing work (Lee et al., 2023) focusing on basic MCL formulation and setting, we expand the formulation and studies to more realistic scenarios and try to answer – how can different models (including Transformers and Mamba) perform in different MCL tasks. Our experiments and analyses show that Mamba can perform well on most of the MCL scenarios. Mamba performs significantly better than other attention-free methods, e.g., Linear Transformers; Mamba can match or outperform the performances of Transformers with fewer parameters and computations. Specifically, on some challenging with more global structures across the sequences (e.g., fine-grained data) and many challenging scenarios (e.g., domain shifts and long sequences), Mamba can perform more reliably and effectively than Transformers, demonstrating better generalization and robustness. Additionally, we analyzed the influence of the model design and conducted preliminary studies to explore the potential of model variants of Mamba, e.g., Mamba mixture-of-experts (MoE), in MCL.
@@ 2 Related Work
Continual learning focuses on mitigating catastrophic forgetting, a significant challenge in model training across sequential tasks (De Lange et al., 2021; Wang et al., 2024). The predominant approaches to continual learning are categorized into three main types: replay-based, regularization-based, and architecture-based methods. Replay-based methods, such as maintaining a memory buffer for old task data, effectively prevent forgetting but are constrained by buffer size and potential privacy issues (Rebuffi et al., 2017; Lopez-Paz & Ranzato, 2017; Chaudhry et al., 2019; Buzzega et al., 2020). Alternatively, generative models can approximate previous data distributions to produce pseudo-samples (Shin et al., 2017; Rostami et al., 2019; Riemer et al., 2019). Regularization-based strategies (Kirkpatrick et al., 2017; Zenke et al., 2017; Nguyen et al., 2017; Li & Hoiem, 2017; Aljundi et al., 2018; Zhang et al., 2020) mitigate forgetting by penalizing changes to critical parameters of previous tasks and employing knowledge distillation to retain earlier knowledge. Lastly, architecture-based methods (Yoon et al., 2017; Serra et al., 2018; Li et al., 2019; Yan et al., 2021; Ye & Bors, 2023) allocate specific subsets of parameters to individual tasks, utilizing techniques like task masking or dynamic architecture adjustment to minimize task interference.
Meta-learning is a learning paradigm where models improve their ability to adapt to new tasks by leveraging limited data and prior experience. The bi-level optimization framework of meta-learning is inherently suited for continual learning, as it focuses on balancing the fit for current tasks while maintaining generalization across all previously encountered tasks (Riemer et al., 2018; Beaulieu et al., 2020; Gupta et al., 2020; Wu et al., 2024). Meta-continual learning (MCL) deviates from traditional continual learning settings by incorporating multiple continual learning episodes, structured into meta-training and meta-testing sets (Son et al., 2023). Lee et al. (2023) conceptualizes MCL as a sequence modeling problem, aligning the continual learning objectives with autoregressive models typical in language modeling. OML (Javed & White, 2019) employs a dual-architecture approach, updating a prediction network while keeping the encoder static during training, then optimizing both components in meta-testing for stability. MetaICL (Min et al., 2022) introduces a meta-training framework for natural language in-context learning. MetaICL sharing a common mathematical formulation with MCL, while the underlying functions to be fitted are distinct. Compared to text sequences, the problems we address are inherently more complex, requiring the learning of more intricate functions and making the learning process more challenging.
Transformer architecture is esteemed for its superior sequence modeling capabilities, largely attributed to its attention mechanism (Vaswani et al., 2017). Decoder-only models like GPT (Brown et al., 2020) and Llama (Touvron et al., 2023), which process inputs causally, have significantly propelled the success of modern deep learning. Although Transformers employing softmax-based attention benefit from efficient parallel training, they encounter challenges due to their quadratic computational complexity relative to sequence length. This has prompted a shift towards more RNN-like models capable of linear-time sequence modeling. As a viable alternative, linear attention substitutes the traditional exponential similarity function with a simple dot product across transformed key/query vectors, gaining traction through recent advancements (Katharopoulos et al., 2020; Choromanski et al., 2020; Tay et al., 2020).
State Space Models (SSMs), inspired by traditional state-space models (Kalman, 1960), have recently emerged as a promising architecture for sequence modeling (Gu et al., 2021a; b). Mamba incorporates time-varying parameters into the SSM framework through a selective architecture and enhances training and inference efficiency with a hardware-aware algorithm (Gu & Dao, 2023; Dao & Gu, 2024). It is widely applied in fields such as computer vision and natural language processing (Zhu et al., 2024; Zhang et al., 2024; Han et al., 2024; Lieber et al., 2024).
@@ 3 Problem Formulation and Methodology
In Continual Learning (CL), given a non-stationary data stream ${\mathcal{D}}_{T}^{\text{train}}=(({\mathbf{x}}_{1},y_{1}),\dots,({\mathbf{x}}_{t},y_{t}),\dots,({\mathbf{x}}_{T},y_{T}))$ as training data, where ${\mathbf{x}}_{t}\in{\mathcal{X}}_{t}$ and $y_{t}\in{\mathcal{Y}}_{t}$. A predictive model $g_{\phi_{t}}():{\mathcal{X}}\rightarrow{\mathcal{Y}}$ is trained on the stream (at step $t$) as $P_{\phi_{t}}(y|{\mathbf{x}})$ for a potential testing set ${\mathcal{D}}_{T}^{\text{test}}=\{({\mathbf{x}}_{n},y_{n})\}_{n=1}^{N}$, where ${\mathbf{x}}_{t}\in{\mathcal{X}}_{t}$ and $y_{t}\in{\mathcal{Y}}_{t}$, with the same distribution as the training set. A conventional continual learner is manually crafted for continual updating/optimizing the model parameter $\phi_{t}$. The data stream ${\mathcal{D}}_{T}$ usually consists of the data from different tasks or distributions, which is usually piecewise stationary within an interval of a task. In a general online CL setting, each sample point can only be seen once; if the samples belonging to one task can be held and accessed as a batch, it is an offline CL. We mainly consider the online CL setting.
@@
For achieving efficient CL, Meta-Continual Learning (MCL) (Lee et al., 2023; Son et al., 2023) is formulated to meta-learn a parameterized continual learner that can efficiently learn/update a predictive model (in ${\mathcal{X}}\rightarrow{\mathcal{Y}}$) from samples in a data stream (in ${\mathcal{X}}\times{\mathcal{Y}}$). Considering that the (continually) learned model from a sequence ${\mathcal{D}}_{t}^{\text{train}}$ is deployed for prediction given a testing sample input ${\mathbf{x}}_{\text{test}}$, i.e., $P_{\theta}(y_{\text{test}}|{\mathbf{x}}_{\text{test}},{\mathcal{D}}_{t}^{\text{train}})$. Thus MCL is equivalent to learning a functional model of the predictive model functions. And MCL can be treated as the task of learning a sequence prediction model $f_{\theta}():({\mathcal{X}}\times{\mathcal{Y}})\times{\mathcal{X}}\rightarrow{\mathcal{Y}}$ parameterized by $\theta$. $f_{\theta}()$ can continually take streaming data ${\mathcal{D}}_{t}^{\text{train}}$ as input and make predictions for any testing samples in a ${\mathcal{D}}^{\text{test}}$ conditioning on ${\mathcal{D}}_{t}^{\text{train}}$ via $\hat{y}_{\text{test}}=f_{\theta}({\mathbf{x}}_{\text{test}},{\mathcal{D}}_{t}^{\text{train}})$. The learner updates internal hidden states to reflect the continually taken data samples corresponding to a CL process. By giving multiple episodes with $({\mathcal{D}}^{\text{train}},{\mathcal{D}}^{\text{test}})$, the parameter $\theta$ of the learner can be learned in the meta-learning/updating process for optimizing the performance on all ${\mathcal{D}}^{\text{test}}$. Note that the targets $y$ in different episodes are independent, which are only symbolic indicators without general semantic meaning across episodes. In this work, we focus on MCL based on a parameterized sequence prediction model for general purpose, despite the existence of other types of meta-learning scheme (Finn et al., 2017; Javed & White, 2019).
3.1 Preliminaries: Transformers, Linear Transformers, and SSMs

### v4

@@ Abstract
Continual learning (CL) aims to efficiently learn from a non-stationary data stream, without storing or recomputing all seen samples. CL enables prediction on new tasks by incorporating sequential training samples. Building on this connection between CL and sequential modeling, meta-continual learning (MCL) aims to meta-learn an efficient continual learner as a sequence prediction model, with advanced sequence models like Transformers being natural choices. However, despite decent performance, Transformers rely on a linearly growing cache to store all past representations, conflicting with CL’s objective of not storing all seen samples and limiting efficiency. In this paper, we focus on meta-learning sequence-prediction-based continual learners without retaining all past representations. While attention-free models with fixed-size hidden states (e.g., Linear Transformers) align with CL’s essential goal and efficiency needs, they have shown limited effectiveness in MCL in previous literature. Given Mamba’s strong sequence modeling performance and attention-free nature, we explore a key question: Can attention-free models like Mamba perform well on MCL? By formulating Mamba and the SSM for MCL tasks, we propose MambaCL, a meta-learned continual learner. To enhance MambaCL’s training, we introduce selectivity regularization, leveraging the connection between Mamba and Transformers to guide its behavior over sequences. Furthermore, we study how Mamba and other models perform across various MCL scenarios through extensive and well-designed experiments. Our results highlight the promising performance and strong generalization of Mamba and attention-free models in MCL, demonstrating its potential for efficient continual learning and adaptation.
@@ 1 Introduction
Continual learning (CL) aims to efficiently learn and accumulate knowledge in a non-stationary data stream containing different tasks, without storing or recomputing training on all previously seen data [11, 64]. Given a sequence of data ${\mathcal{D}}_{T}=(({\mathbf{x}}_{1},y_{1}),...,({\mathbf{x}}_{t},y_{t}),...,({\mathbf{x}}_{T},y_{T}))$ with a series of paired observations ${\mathbf{x}}_{i}$ (e.g., images) and targets $y_{i}$ (e.g., class labels) from different tasks, CL is usually produced to learn one model $P_{\phi_{t}}(y|{\mathbf{x}})$ parameterized by $\phi_{t}$ that can perform prediction for any tasks corresponding to the seen data ${\mathcal{D}}_{t}$. For example, in class incremental learning (CIL) [49, 73], a widely studied CL scenario, ${\mathcal{D}}_{T}$ consists of data with incrementally added classes, and $P_{\phi_{t}}(y|{\mathbf{x}})$ is trained to recognize all seen classes. To ensure efficiency, CL methods learn from data streams while minimizing historical data storage and constraining memory growth, ideally at least a sub-linear rate [11, 44, 63]. The key challenge in CL is maintaining performance on past tasks while continuously updating model parameters $\phi_{t}$ [11, 64].
@@
CL methods continually train/update the model $P_{\phi_{t}}(y|{\mathbf{x}})$ from seen sequence ${\mathcal{D}}_{t}$ at arbitrary step $t$, enabling prediction of the output $y^{\text{test}}$ for any observation ${\mathbf{x}}^{\text{test}}$. The whole learning and prediction process can be viewed as a sequence prediction (SP) problem [32, 3], where the goal is to predict $y_{\text{test}}$ for a query ${\mathbf{x}}_{\text{test}}$ conditioning on the seen data sequence and the testing input, i.e., $({\mathcal{D}}_{t}^{\text{train}},{\mathbf{x}}^{\text{test}})\equiv({\mathbf{x}}_{1}^{\text{train}},y_{1}^{\text{train}},...,{\mathbf{x}}_{t}^{\text{train}},y_{t}^{\text{train}},{\mathbf{x}}^{\text{test}})$. In conventional CL, the model parameter $\phi_{t}$ is trained to maintain the historical knowledge (as a states on the sequence) through an optimization step $\phi_{t+1}=\texttt{optim-step}(\phi_{t},{\mathbf{x}}_{t},y_{t})$. This connection between SP and CL motivates meta-learning a continual learner as a sequence prediction model, for computation-and-data-efficient CL. Through meta-continual learning (MCL) framework [32, 58, 3], a continual learner $f_{\theta}()$ parameterized by $\theta$ can be trained multiple CL episodes. A meta-learned $f_{\theta}()$ can take a sequence $({\mathcal{D}}_{t},{\mathbf{x}}_{t+1})$ as input and predict $y_{t+1}=f_{\theta}(({\mathcal{D}}_{t},{\mathbf{x}}_{t+1}))$, equivalent to a predictive model $P_{\theta}(y|{\mathbf{x}},{\mathcal{D}}_{t})$. The data stream serves as a context for making predictions on new queries.
Considering the strong sequence modeling ability, Transformers [65, 61] are a natural choice for SP-based MCL [32, 58]. In MCL, given a data stream, a meta-learned Transformer generates a new key-value pair at each step and maintains the key-value pairs for all seen training samples in a key-value cache. The prediction for the query at each step is based on attention over the key-value pairs retained from all preceding training samples. This retrieval-based approach enables strong performance in CL and MCL datasets [32, 3]. Benefiting from the retrieval-based modeling and explicit intra-sequence attention, Transformers can perform effectively on the datasets formulated for CL and MCL [32, 3]. However, Transformers rely on maintaining the key-value cache of all seen samples. While the key-value cache resembles a recurrent neural network (RNN) hidden state [27, 32], its size grows linearly with seen tokens, increasing memory and computation costs, which contradicts CL’s goal of learning without storing full historical data. The basic Transformer may not fully fit the intention of MCL, despite the reported decent performance [32].
In this paper, we focus on MCL without retaining all past representations. Among various approaches to improving Transformer efficiency [60], attention-free models achieve efficiency by approximating softmax attention with kernel methods and linear operations, resulting in constant hidden state sizes and linear computation complexity, e.g., Linear Transformer (Linear TF) [27] and Performer [8]. While these models align better with CL’s goals, they struggle in MCL due to limited expressive power [32, 27, 8, 60]. Recent advances in state-space models (SSMs) have led to efficient attention-free models for long-sequence processing with nearly linear computation [17, 18]. By incorporating time-varying modeling, Mamba [16, 9] achieves outstanding performance in language modeling [15].
@@
Given Mamba’s strong performance as an attention-free model, we ask: Can the attention-free model like Mamba perform well in MCL? To explore this, we propose MambaCL, formulating selective SSM for MCL and meta-learning Mamba as an online continual learner. Training SP models for MCL proves challenging, especially for Mamba, which lacks explicit attention operations as in Transformers. To address this, we introduce a selectivity regularizer leveraging connections between SSM/Mamba, Linear Transformers, and Transformers for training, guiding the learning of time-variant operations and behaviors in the selective SSM during training. Beyond the scope of the existing works [32] focusing on basic MCL formulation and setting, we expand the studies to more realistic scenarios and try to answer: how different models (including Transformers and Mamba) perform across various MCL scenarios. Our experiments show that Mamba can perform well in most MCL scenarios, significantly outperforming other attention-free methods like Linear Transformers and matching or surpassing Transformers with fewer parameters and computations. Notably, in some challenging settings with strong global structures (e.g., fine-grained) and scenarios with domain shifts or long sequences, Mamba demonstrates promising reliability, generalization, and robustness, highlighting its potential for efficient CL and adaptation. We also analyze model design influences and explore potential Mamba variants, such as incorporating mixture-of-experts for MCL.
@@ 2 Related Work
Continual learning aims to address catastrophic forgetting when training models across sequential tasks [11, 64]. Existing approaches include replay, regularization, and architecture-based methods. Replay-based techniques retain task knowledge via memory buffers [49, 37, 7, 6] or generate pseudo-samples using generative models [57, 52, 51], though they face memory and privacy constraints. Regularization-based methods preserve previous knowledge by penalizing important parameter changes or applying distillation [28, 70, 43, 34, 1, 71]. Architecture-based methods mitigate interference by assigning task-specific parameters via masking or dynamic adaptation [69, 56, 33, 67, 68, 63, 38].
Meta-learning aims to learn generalizable models across tasks, enabling efficient adaptation to new tasks with limited data. The meta-learning is inherently suited for and has been adapted for CL, focusing on balancing the fit for current tasks while maintaining generalization across all previously encountered tasks [50, 2, 21, 66, 41, 42, 55, 20, 4, 39, 23]. Meta-continual learning (MCL) deviates from traditional CL settings by incorporating multiple CL episodes, structured into meta-training and meta-testing sets [58, 32, 3]. [32] frames MCL as a sequence modeling task, aligning continual learning with language modeling. OML [25] adopts a dual-architecture, updating predictor while freezing encoder during training, then jointly optimizing both during meta-testing for stability.
Transformer architectures excel in sequence modeling due to their attention mechanism [65]. Decoder-only variants like GPT [5] and LLaMA [61], which process inputs causally, have driven major advances in deep learning. However, their softmax-based attention incurs quadratic complexity with sequence length, limiting scalability. This has spurred interest in RNN-like alternatives with linear-time modeling. In particular, linear attention methods replace the exponential similarity with dot products over transformed key/query vectors, offering improved efficiency [27, 8, 59].
State Space Models (SSMs), inspired by classical control theory [26], have recently gained traction for sequence modeling [17, 18]. Mamba advances this framework by introducing time-varying parameters and a hardware-efficient implementation [16, 9], achieving fast training and inference. It has proven effective in vision and language tasks [74, 72, 22, 35].
@@ 3 Problem Formulation and Methodology
In Continual Learning (CL), given a non-stationary data stream ${\mathcal{D}}_{T}^{\text{train}}=\{({\mathbf{x}}_{t},y_{t})\}_{t=1}^{T}$ with ${\mathbf{x}}_{t}\in{\mathcal{X}}_{t}$ and $y_{t}\in{\mathcal{Y}}_{t}$ as training data, a predictive model $g_{\phi_{t}}():{\mathcal{X}}\rightarrow{\mathcal{Y}}$ is trained on the stream as $P_{\phi_{t}}(y|{\mathbf{x}})$ (at step $t$) for a potential testing set ${\mathcal{D}}_{T}^{\text{test}}=\{({\mathbf{x}}_{n},y_{n})\}_{n=1}^{N}$, where ${\mathbf{x}}_{t}\in{\mathcal{X}}_{t}$ and $y_{t}\in{\mathcal{Y}}_{t}$, following the distribution of the training set. A conventional continual learner is manually crafted to continually update model parameters $\phi_{t}$ through optimization. The data stream ${\mathcal{D}}_{T}$ typically comprises samples from different tasks or distributions, often piecewise stationary within a task interval. In online CL, each sample is seen only once, whereas in offline CL, task-specific samples can be accessed for multiple rounds. This work primarily focuses on the online CL setting.
@@
For achieving efficient CL, Meta-Continual Learning (MCL) [32, 58] meta-learns a parameterized continual learner that updates a predictive model (in ${\mathcal{X}}\rightarrow{\mathcal{Y}}$) from a streaming data source (in ${\mathcal{X}}\times{\mathcal{Y}}$). Given a learned model from a sequence ${\mathcal{D}}_{t}^{\text{train}}$ in MCL, prediction can be formulated as $P_{\theta}(y_{\text{test}}|{\mathbf{x}}_{\text{test}},{\mathcal{D}}_{t}^{\text{train}})$. MCL can be seen as training a sequence prediction (SP) model $f_{\theta}():({\mathcal{X}}\times{\mathcal{Y}})\times{\mathcal{X}}\rightarrow{\mathcal{Y}}$, where $f_{\theta}()$ continually processes streaming data and performs prediction for test samples via $\hat{y}_{\text{test}}=f_{\theta}({\mathbf{x}}_{\text{test}},{\mathcal{D}}_{t}^{\text{train}})$. The learner updates its hidden states dynamically, reflecting the continual learning process. Through multiple episodes of $({\mathcal{D}}^{\text{train}},{\mathcal{D}}^{\text{test}})$, the learner’s parameters $\theta$ are optimized via meta-learning to improve performance across all. The targets $y$ in different episodes are independent, not sharing semantic meaning. This work focuses on MCL using a parameterized SP model, despite other meta-learning approaches [14, 25].
3.1 Preliminaries: Transformers and SSMs